OpenStudy (anonymous):

the are of a square is 49 sq units. find the length of the side of the square algebraically.

OpenStudy (anonymous):

area=length^2

OpenStudy (anonymous):

length=area^1/2

OpenStudy (anonymous):

I don't get it :(

OpenStudy (anonymous):

area=49 let the length be x area=length*length area=x*x x^2=49 x=7,-7 since it cant be -7 x=7

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

we knw tht.................ar of square = side x side ...nn area is given = 49 so....side= 7